A box contains 15 large marbles and 19 small marbles. Each marble is either green or white. 7 of the large marbles are green, and 6 of the small marbles are white. If a marble is randomly selected from the box, what is the probability that it is small or white? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

Total marbles = 15 + 19 = 34

P(Marble is small or white)

= P(Small) + P(White) - P(Small white)

= 19/34 + (8 + 6)/34 - 6/34

= 27/34

= 0.7941